Trying to figure out an equation for the following:
Create a model for how old in years each of you and the oldest person you chose will be when your ages add up to 150 years old. The numbers are 33 and 57.
For example, if x = your age and the eldest person was a year older than you, you would write their age as x + 1.

Answer by Theo(13342)

if you're 33 and the other person is 57, then your ages add up 90.
150 - 90 = 60.
if you let x equal the number of years each of you have to age, then you get 2x = 60, and when you solve for x, you get x = 30.
you and the other person's age will be equal to 150 in 30 years.
33 + 30 = 63.
57 + 30 = 87.
63 + 87 = 150.

if you wish to generalize this, then you might do something like the following.

your age is x.
the oldest person you know is x + a, where a is the difference in age between you and the other person.

if you let b equal the number of years each of you have to age, then your formula becomes:

2b = 150 - 2x - a.

solving for b, the formula becomes b = (150 - 2x - a) / 2.

if you're 33, then x = 33.
if your oldest adult is 57, then a = 57 - 33 = 24.
formula becomes b = (150 - 2 * 33 - 24) / 2 which becomes b = (150 - 90) / 2 which becomes b = 60 / 2 which becomes b = 30.

this formula should work for any age.
it will be possible to get partial years, since you're not guaranteed you will always get a number that is divisible by 2 resulting in an integer.

the formula is b = (150 - 2x - a) / 2.

x is your age
a is the difference between your age and the age of the oldest person you know.
b is how many years each of you would have to add to your age to make the sum equal to 150.

there may be a simpler way to do it.

let x = your age.
let y = the oldest person you know age.
b = how much each of you have to add to your ages.

the formula would be b = (150 - (x+y)) / 2

all you have to do is add up your ages, subtract them from 150, and divide by 2.

assuming you are 21 and the oldest adult you know is 72.
add them up and you get 93.
subtract that from 150 and you get 57.
divide that by 2 and you get 28.5 years each.
21 + 28.5 = 49.5
72 + 28.5 = 100.5
100.5 + 49.5 = 150

the formula is b = (150 - (x+y))/2
x = 21
y = 72
formula becomes b = (150 - 93)/2 which becomes 57/2 which becomes 28.5.

you can do it either way.